OpenStudy (anonymous):

5(3x-4)+4x=4(5x-2)-12 Use additive property of equality to solve for x. Please show step by step so I know If I'm doing correctly

OpenStudy (anonymous):

1) step one is distribute on both sides 2) combine like terms on both sides 3) subtract \(19x\) from both sides 4) last step...

OpenStudy (anonymous):

Huh! I'm sorry but this is all new to me I would have to actually see how you did it. I got it to 20 x so I must be doing something wrong!

OpenStudy (anonymous):

\[5(3x-4)+4x=4(5x-2)-12\]1) distribute gives \[15x-20+4x=20x-8-12\]

OpenStudy (anonymous):

2) combining like terms gives \[19x-20=20x-20\]

OpenStudy (anonymous):

2) subtract \(19x\) \[-20=x-20\] add \(20\) \[0=x\]

OpenStudy (anonymous):

Thank you. Stupid mistake. I forgot to multiply 5 with 3 x and just put down 3x! It totally changed the problem. Thank you so much!